OpenStudy (anonymous):

the dollar profit from the expenditure of x thousand dollars on advertising is given by p(x)=780+35x-2x^2. find the marginal profit when the expenditure is x=14.

OpenStudy (anonymous):

Simply put x = 14 into your function P(x) Your profit P(x) = 780 + 35x - 2x², so when x = 14 your profit will be: P(14) = 780 + 35(14) - 2(14²) = $878

OpenStudy (anonymous):

not an answer choice. i got it already